OpenStudy (anonymous):

Use the addition-subtraction method to find all solutions of the system of equations. Suggestion: first clear both equations of fractions. 1/4s-1/3t=4 2/7s-1/7t=1/10

jhonyy9 (jhonyy9):

3s-4t=48 20s-10t=7

jhonyy9 (jhonyy9):

multiply first by -20 and second by 3 -60s+80t=-960 60s-30t=212 --------------- 0 50t= - 748 t= -748/50 = -374/25 20s- 10(-374/25)=7 20s+748/5=7 100s= - 713 s= -713/100